### v1

# HyGEN: Regularizing Negative Hyperedge Generation for Accurate Hyperedge Prediction

Abstract.
Hyperedge prediction is a fundamental task to predict future high-order relations based on the observed network structure. Existing hyperedge prediction methods, however, suffer from the data sparsity problem. To alleviate this problem, negative sampling methods can be used, which leverage non-existing hyperedges as contrastive information for model training. However, the following important challenges have been rarely studied: (C1) lack of guidance for generating negatives and (C2) possibility of producing false negatives. To address them, we propose a novel hyperedge prediction method, HyGEN, that employs (1) a negative hyperedge generator that employs positive hyperedges as a guidance to generate more realistic ones and (2) a regularization term that prevents the generated hyperedges from being false negatives. Extensive experiments on six real-world hypergraphs reveal that HyGEN consistently outperforms four state-of-the-art hyperedge prediction methods.
Hyperdge prediction, Adversarial learning, Negative sampling
journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3715456isbn: 979-8-4007-1331-6/25/04ccs: Computing methodologies Adversarial learningccs: Computing methodologies Regularizationccs: Computing methodologies Network science
1. Introduction
A hypergraph, a generalized data structure, is capable of modeling a high-order (i.e., group-wise) relation among an arbitrary number of real-world objects as a hyperedge. Due to its powerful expressiveness, hypergraph-based network learning (et al., 2019a, 2022b) has been widely studied and shown to outperform graph-based methods in various downstream tasks, including node classification (et al., 2020b, 2023a), node ranking (et al., 2021c, b), and link prediction (et al., 2020e; Vaida and Purcell, 2019).
Figure 1. Process of hyperedge prediction.
Hyperedge prediction (i.e., link prediction on hypergraphs) is a fundamental task in many real-world applications, such as recommender systems (et al., 2021c) and social network analysis (et al., 2021b, 2023a); it predicts future high-order relations (i.e., hyperedges) based on an observed hypergraph structure. A general approach to hyperedge prediction is two-fold (et al., 2022a, 2023b) (see Figure 1: (1) (hypergraph encoding) the embeddings of nodes are produced by hypergraph neural networks (et al., 2019a, 2020b) and (2) (hyperedge candidate scoring) the embeddings of nodes in each hyperedge candidate are aggregated and fed into a predictor to determine whether the candidate is real.
In real-world networks, however, high-order relations are often extremely sparse (et al., 2020c) (i.e., $2^{|V|}\gg|E|$, where $V$ and $E$ are the sets of nodes and hyperedges, respectively). Such a data sparsity problem is the fundamental cause of low accuracy in hyperedge prediction. To address this problem, negative sampling (NS) can be used (et al., 2020e, 2019b, 2022a), utilizing non-existing hyperedges as contrastive information for model training. Specifically, the model is trained so that positive examples get higher scores while negative examples get lower scores, which enhances the distinguishing ability of the model. Thus, it is crucial to carefully choose negative hyperedges to maximize the effect of negative sampling.
However, sampling ‘good’ negative hyperedges is challenging in the context of hyperedge prediction since there exist too many possible negative hyperedges (i.e., $2^{|V|}-|E|$). Although existing methods (et al., 2019b, 2020e), enhanced by NS, have achieved breakthroughs in many fields, they focus primarily on hypergraph encoding while employing simple heuristic-based NS methods (et al., 2020c). Thus, negative sampling for hyperedge prediction is still rarely explored.
Although one recent work (et al., 2022a) proposed an adversarial-training-based hyperedge prediction method (AHP) that leverages model-generated negative hyperedges useful for model training; however, it has overlooked the following two important challenges:
(C1) Lack of guidance for generating negatives. Only a random noise is used as the input of the hyperedge generator in AHP. Thus, it may fail to effectively reflect the characteristics of positive hyperedges into generated hyperedges, especially in the initial stage of the training, which could lead to inefficient and unstable training.
(C2) Possibility of producing false negatives. AHP generates negative hyperedges by using a generative adversarial network (GAN), which aims to generate negative hyperedges similar as much as possible to positive hyperedges (i.e., copying the exact distribution of the original hyperedges). Without any regularization, however, it might be possible to generate hyperedges too similar to positives, which would be potential positive hyperedges.
From this motivation, in this paper, we propose a novel adversarial-training-based method, HyGEN which stands for regularizing negative Hyperedge GENeration for accurate hyperedge prediction. HyGEN employs (1) a positive-guided negative hyperedge generator that leverages positive hyperedges as guidance to generate more-realistic negative hyperedges for (C1) and (2) a regularization term to prevent the generated hyperedges from being too similar to positive hyperedges for (C2).
Contributions. The main contributions of this work are as follows.
  • Challenges: We point out two important challenges of negative sampling in hyperedge prediction: (C1) lack of guidance for generating negatives and (C2) possibility of false negatives.
  • Method: We propose a novel hyperedge prediction method, HyGEN that employs (1) a positive-guided negative hyperedge generator for (C1) and (2) a regularization term for (C2).
  • Evaluation: Via extensive experiments on six real-world hypergraphs, we verify the superiority of HyGEN over four state-of-the-art hyperedge prediction methods.
2. Related Works
Hyperedge prediction. There have been a number of works to study hyperedge prediction. Expansion (et al., 2020d) models a hypergraph as multiple n-projected graphs and predicts future hyperedges based on the projected graphs. HyperSAGNN (et al., 2019b) uses a self-attention-based GNN model to learn hyperedges of variable sizes. NHP (et al., 2020e) employs hyperedge-aware GNN models to learn node embeddings in hypergraphs, using the max-min pooling to aggregate the embeddings of nodes within each hyperedge candidate for prediction. AHP (et al., 2022a), the state-of-the-art hyperedge prediction method, employs adversarial training to generate negative hyperedges for model training and uses max-min pooling for node aggregation.
Negative hyperedge sampling. For enhancing the training of hyperedge prediction models, the following three heuristic-based methods for negative hyperedge sampling have been proposed (et al., 2020c): (1) Sized NS (SNS) samples $n$ nodes uniformly at random; (2) Motif NS (MNS) transforms a hypergraph into an ordinary graph via a clique-expansion and samples a $n$-connected component in the expanded graph; and (3) Clique NS (CNS) selects a hyperedge $e$ and replaces one of its incident nodes $u\in e$ with a node $v\notin e$, which is linked to all the other incident nodes, i.e., ($e\setminus\{u\})\cup\{v\}$.
3. Proposed Method: HyGEN
Notations. A hypergraph is defined as $H=(V,E)$, where $V=\{v_{1},v_{2},\dots,v_{|V|}\}$ and $E=\{e_{1},e_{2},\dots,e_{|E|}\}$. A hypergraph can generally be represented by an incidence matrix $\mathbf{H}\in\{0,1\}^{|V|\times|E|}$, where each element $h_{ij}=1$ if $v_{i}\in e_{j}$, and $h_{ij}=0$ otherwise. The node and hyperedge features are represented by the matrices $\mathbf{P}\in\mathbb{R}^{|V|\times d}$, $\mathbf{Q}\in\mathbb{R}^{|E|\times d}$, where each row $p_{i}$ and $q_{i}$ represents the $d$-dimensional feature of a node and a hyperedge, respectively.
Problem 1 (Hyperedge Prediction). Given a hypergraph $\mathbf{H}\in\{0,1\}^{|V|\times|E|}$ and the initial node features $\mathbf{X}\in\mathbb{R}^{|V|\times d}$, and a hyperedge candidate $e^{\prime}\notin E$, to predict whether $e^{\prime}$ is real or not.
Figure 2. Overview of HyGEN: (1) hypergraph encoding (upper) and (2) hyperedge candidate scoring (lower).
Overview of HyGEN. Figure 2 illustrates the overview of HyGEN, which consists of (1) hypergraph encoding (upper) and (2) hyperedge candidate scoring (lower) that we focus on.
(1) Hypergraph encoding. Given a hypergraph $H=(V,E)$, HyGEN produces node embeddings $\mathbf{P}\in\mathbb{R}^{|V|\times d}$ and hyperedge embeddings $\mathbf{Q}\in\mathbb{R}^{|E|\times d}$. Following (et al., 2020b, 2021a), HyGEN adopts a 2-stage aggregation approach, which repeats (1) (node-to-hyperedge) producing a hyperedge embedding by aggregating the node embeddings and (2) (hyperedge-to-node) producing a node embedding by aggregating the hyperedge embeddings. Formally, the node and hyperedge embeddings at the l-th layer are defined as:
(1) $\footnotesize\mathbf{Q}^{(l)}=\sigma\left(\mathbf{H}^{T}\mathbf{P}^{(l-1)}\mathbf{W}^{(l)}_{E}+b^{(l)}_{E}\right),\footnotesize\mathbf{P}^{(l)}=\sigma\left(\mathbf{H}\mathbf{Q}^{(l)}\mathbf{W}^{(l)}_{V}+b^{(l)}_{V}\right),$
where $\mathbf{P}^{(0)}=\mathbf{X}$, $\mathbf{W}^{(l)}_{*}$ and $b^{(l)}_{*}$ are trainable weight and bias matrices, respectively; $\sigma$ is a non-linear activation function; normalization terms are omitted for simplicity in Eq. 1.
(2) Hyperedge candidate scoring. The hyperedge candidate scoring of HyGEN consists of a (a) generator to produce informative negative hyperedges for training and a (b) discriminator to predict whether a hyperedge candidate is positive or negative.
(2)-(a) Generator. We propose a positive-guided negative hyperedge generator for (C1) that employs an encoder-decoder structure. Specifically, given a positive hyperedge $e^{+}$, (1) the encoder produces the latent vector $q_{e^{+}}$ for positive hyperedge $e^{+}$ and (2) the latent vector $q_{e^{+}}$ and a random Gaussian noise $z$ are fed into the decoder to generate a node membership vector $c\in\mathbb{R}^{|V|}$:
[…]
where $e^{+}\in\mathbb{R}^{|V|}$ is a one-hot vector whose element $e^{+}[i]=1$ if $i\in e^{+}$, and $e^{+}[i]=0$ otherwise, and $c_{i}$ is a node membership vector whose element $c_{i}$ represents the probability of the node $i$ being included in the generated negative hyperedge.
For effectively extracting the characteristics of positive hyperedges and injecting them into generated negative hyperedges, inspired by (et al., 2020a), we adopt a convolutional neural network (CNN) with three layers as the architecture of our encoder and decoder. Each of the layers consists of a 1-D convolutional layer with 256 kernels of size 3, average-pooling, and LeakyReLU as the activation function. In the case of the decoder, we additionally use adaptive instance normalization (AdaIN) (et al., 2017b), which follows each convolutional layer, in order to inject the characteristics of a positive hyperedge into its corresponding negative hyperedge generated.
Finally, for the size $n$ of a negative hyperedge, sampled from the size distribution of positive hyperedges, HyGEN selects top-$n$ nodes as a negative hyperedge $e^{-}$ from the candidate probability vector $c$. As a result, the pair of positive and negative hyperedges $e^{+}$ and $e^{-}$ are fed into the discriminator for model training.
(2)-(b) Discriminator. Given the learned node embeddings $\mathbf{P}$ and a hyperedge candidate $e^{\prime}$ ($e^{+}$ or $e^{-}$), the discriminator (1) produces the embedding of a hyperedge candidate $q_{e^{\prime}}$ by aggregating the embeddings of the nodes in $e^{\prime}$, $\mathbf{P}[e^{\prime},:]\in\mathbb{R}^{|e^{\prime}|\times d}$, and (2) computes the probability $\hat{y}_{e^{\prime}}$ of $e^{\prime}$ being formed based on $q_{e^{\prime}}$ as:
(3) $agg(\mathbf{P}[e^{\prime},:])\rightarrow q_{e^{\prime}}\in\mathbb{R}^{|d|},\hskip 8.53581ptpred(q_{e^{\prime}})\rightarrow\hat{y}_{e^{\prime}}\in\mathbb{R}^{1},$
where $agg(\cdot)$ is the element-wise maxmin pooling, used as the aggregation function to reflect the diversity of the embeddings of nodes in a hyperedge candidate, by following (et al., 2022a, 2020e), and $pred(\cdot)$ is a hyperedge predictor, which consists of three fully-connected layers ($d\times 128\times 8\times 1$), followed by a sigmoid function.
(3) Model training. We train the model parameters of HyGEN in an adversarial way (et al., 2017a): given a batch $B$ of positive hyperedges, (1) generate $|B|$ negative hyperedges using the generator $G$ ($enc(\cdot)$ and $dec(\cdot)$), (2) classify the positive and negative hyperedges using the discriminator $D$ ($agg(\cdot)$ and $pred(\cdot)$), and (3) update the model parameters of HyGEN based on their losses. Specifically, as $D$ aims to compute the probabilities of positive hyperedges higher than those of negative hyperedges, the loss function for $D$ is defined as:
(4) $L_{D}=-\frac{1}{|B|}\sum_{e^{+}\in B}[D(e^{+}|H,X)]+\frac{1}{|B|}\sum^{|B|}_{j=1}[D(G(z_{j}\mid e^{+}_{j})|H,X)],$
where $e^{+}$ is a positive hyperedge and $G(z_{j}\mid e^{+}_{j})$ is the negative hyperedge generated from a noise $z$ and the positive hyperedge $e^{+}_{j}$. This loss $L_{D}$ is also used for training the hypergraph encoder $f(\cdot)$.
On the other hand, $G$ aims to deceive $D$ to misclassify negative hyperedges as positive. Thus, the loss function for $G$ is defined as:
(5) $L_{G}=-\frac{1}{|B|}\sum^{|B|}_{j=1}[D(G(z_{j}\mid e^{+}_{j})|H,X)].$
Regularization for (C2). ‘Hard’ negative hyperedges, generated by our hyperedge generator, could enhance the distinguishing ability of a hyperedge prediction model (et al., 2022a). Without any regularization on a generator, however, it may completely copy the distribution of the original positive hyperedges; result in generating hyperedges too similar to positive hyperedges that might potentially become positive hyperedges in the future. Using such hyperedges as negative hyperedges in training can lead to incorrect learning.
To address this challenge, we propose a novel regularization term that is integrated into the loss function. This regularization loss gives a smaller penalty as the generated negative hyperedges are more similar to positive hyperedges until a certain degree and gives a larger penalty as they are too similar to positive hyperedges. Specifically, given the embeddings of positive and negative hyperedges $\mathbf{Q}^{+}$ and $\mathbf{Q}^{-}$, the regularization loss is defined as:
(6) $L_{reg}=-\left(\frac{|\theta-k|}{\theta(1-\theta)}\right)^{p},\hskip 8.53581pt\theta=sim(\mathbf{Q}^{+},\mathbf{Q}^{-}).$
where $k$ and $p$ are hyperparameters that control the converge point and the curvature of the function, respectively (See Figure 3). $sim(\cdot)$ is the cosine similarity used as a similarity function. Finally, the regularization loss is integrated into the total loss with a hyperparameter $\beta$ to control the weight of the regularization loss:
(7) $L_{total}$ $=L_{D}+L_{G}+\beta\cdot L_{reg}$
Figure 3. Regularization loss according to hyperparameters.
4. Experimental Validation
We evaluate HyGEN by answering the following questions.
  • EQ1 (Accuracy). To what extent does HyGEN improve the existing hyperedge prediction methods in terms of the accuracy?
  • EQ2 (Ablation study). Is each of our strategies beneficial to generating negative hyperedges useful for model training?
  • EQ3 (Sensitivity). How sensitive is the effect of the regularization loss in model training to its hyperparameters ($k$ and $p$)?
4.1. Experimental Setups
Datasets and competitors111 We have released all the code and datasets at: https://github.com/ssong915/HyGEN. We use six widely used real-world hypergraphs: (1) three co-citation datasets (Citeseer, Cora, and Pubmed), (2) two authorship datasets (Cora-A and DBLP-A), and (3) one collaboration dataset (DBLP). We select four state-of-the-art hyperedge prediction methods as our competitors (Expansion (et al., 2020d), NHP (et al., 2020e), HyperSAGNN (et al., 2019b), and AHP (et al., 2022a)).
Evaluation protocol. We evaluate HyGEN by using the protocol exactly same as that used in (et al., 2022a). For each dataset, we use five data splits, where positive hyperedges are randomly divided into training (60%), validation (20%), and test (20%) sets. We use three validation and test sets constructed with negative hyperedges sampled by SNS, MNS, and CNS, explained in Section 2. As metrics, we use AUROC (area under the ROC curve) and AP (average precision). We (1) measure AUROC and AP on each test set when the averaged AUROC over the validation sets is maximized, and (2) report the averaged AUROC and AP over five runs on each test set.
Table 1. Hyperedge prediction accuracy: HyGEN always outperforms all competing methods in both the average AP and AUROC across all datasets. (The bold and underlined represents the best and the second-best results in each test set, respectively.)
AUROC (%) AP (%) AUROC (%) AP (%) AUROC (%) AP (%)
SNS MNS CNS AVG SNS MNS CNS AVG SNS MNS CNS AVG SNS MNS CNS AVG SNS MNS CNS AVG SNS MNS CNS AVG
Expansion Citeseer 66.3 78.1 33.1 59.2 76.5 81.7 49.8 69.3 Pubmed 52.0 73.0 24.1 49.7 67.5 75.5 44.0 62.3 Cora 47.0 70.7 25.6 47.8 63.7 76.4 45.4 61.8
NHP 99.1 70.1 51.0 73.4 99.0 73.1 52.0 74.7 97.3 69.4 52.4 73.0 97.3 65.6 51.3 71.4 94.3 64.1 47.2 68.5 94.9 67.8 50.9 71.2
HyperSAGNN 54.0 41.0 47.3 47.4 62.7 45.5 49.7 52.6 52.5 68.6 54.6 58.6 53.4 68.0 52.9 58.1 61.7 52.7 49.4 54.6 68.7 57.4 50.8 59.0
AHP 94.3 88.1 65.1 82.5 95.2 87.0 66.0 82.7 91.7 84.0 53.3 76.3 91.8 83.4 52.6 75.9 96.4 86.0 57.2 79.9 96.1 83.7 55.2 78.3
HyGEN 98.4 92.6 67.6 86.2 98.5 91.2 69.4 86.4 92.1 87.1 51.6 77.0 93.2 89.0 55.1 79.1 99.1 90.7 58.4 82.7 99.0 89.6 57.0 81.9
Expansion Cora-A 69.0 84.2 43.4 65.5 69.0 87.6 57.7 71.4 DBLP 64.5 80.1 36.6 60.4 75.1 85.6 51.8 70.8 DBLP-A 63.4 82.6 35.0 60.3 73.0 85.2 51.2 69.8
NHP 90.9 67.2 55.0 71.0 92.5 72.0 58.5 74.3 66.3 54.0 50.3 56.9 60.8 52.3 50.1 54.4 96.6 62.3 55.5 71.5 96.5 60.4 53.4 70.1
HyperSAGNN 38.6 59.1 54.2 50.6 53.2 64.3 54.5 57.3 44.8 57.4 57.2 53.1 56.2 60.2 58.6 58.3 54.8 79.1 56.3 63.4 68.6 80.5 55.2 68.1
AHP 95.8 92.4 78.2 88.8 95.7 89.8 79.6 88.4 94.6 82.0 56.8 77.8 94.7 81.5 56.1 77.4 91.6 92.6 66.8 83.7 92.8 92.8 70.7 85.4
HyGEN 97.7 94.7 80.3 90.9 97.7 91.4 82.9 90.7 95.8 83.9 59.1 79.6 95.8 82.6 57.3 78.6 91.3 92.7 67.1 83.7 92.7 92.9 71.5 85.7
Table 2. Ablation study: each of our strategies is always beneficial to improving the accuracy of HyGEN.
Method Citeseer Pubmed Cora Cora-A DBLP DBLP-A
AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP AUROC AP
HyGEN 86.2 86.4 77.0 79.2 82.7 81.9 90.9 90.7 79.6 78.6 83.7 85.7
w/o positive-guided Gen. 85.0 85.0 76.9 79.1 80.9 80.9 90.2 90.0 76.3 76.5 83.0 83.2
w/o $L_{reg}$ 82.4 82.9 76.8 77.1 78.5 78.2 89.5 89.4 72.7 72.4 70.9 61.6
4.2. Experimental Results
EQ1. Hyperedge prediction accuracy. We compare HyGEN with four competing methods in the hyperedge prediction task. Table 1 shows that HyGEN consistently outperforms all competing methods in terms of both the averaged AP and AUROC over three test sets across all datasets. We note that these improvements of HyGEN over AHP (the best competitor) are remarkable, given that AHP has already improved other existing methods significantly in those datasets. Via the t-tests with a 95% confidence level, we verify that the improvements of HyGEN are statistically significant (i.e., the p-values $\leq 0.05$). As a result, these results demonstrate that HyGEN can generate informative negative hyperedges by effectively addressing the two challenges of negative sampling, thereby enhancing the accuracy of a hyperedge prediction task.
Figure 4. The hyperparameter sensitivity of HyGEN.
EQ2. Ablation study. We verify the effectiveness of our proposed strategies individually by ablating one of them: (i) positive-guided negative hyperedge generator and (ii) regularization term. Table  2 shows that the original version of HyGEN always achieves the highest accuracy across all datasets, which indicates that ablating one of our proposed strategies could lead to accuracy degradation. These results verify the effectiveness of our proposed strategies for addressing the two challenges: (C1) lack of guidance for generating negatives and (C2) possibility of false negatives.
EQ3. Sensitivity analysis. We evaluate the impact of hyperparameters $k$ and $p$ on the accuracy of HyGEN. We measure the model accuracy of HyGEN with varying $k$ from 0 to 1.0 in step of 0.1 and $p$ from 1 to 5 in step of 1. As shown in Figure 4, where the $x$-axis represents the converge point hyperparameter $k$, the $y$-axis represents the curvature hyperparameter $p$, and the $z$-axis represents the averaged AUROC, HyGEN with $k\geq 0.4$ consistently achieves higher accuracy than HyGEN with $k<0.4$ regardless of $p$ (i.e., the red wide area on the surface in Figure 4). Based on these results, we believe that the accuracy of HyGEN is insensitive to the regularization hyperparameters $k$ and $p$.
5. Conclusion
In this paper, we identified two key challenges of negative hyperedge sampling in the hyperedge prediction task: (C1) lack of guidance for generating negatives and (C2) possibility of producing false negatives. To address both challenges, we propose a novel hyperedge prediction method, HyGEN that employs (1) a positive-guided negative hyperedge generator leveraging positive hyperedges as a guidance to generate informative negative hyperedges for (C1) and (2) a regularization term to prevent the generated hyperedges from being false negatives (C2). Comprehensive experiments on six real-world datasets verified the superiority of HyGEN over four state-of-the-art hyperedge prediction methods.
Acknowledgements.
This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2022-00155586, 2022-0-00352).